# Solar Cruiser Disturbance Torque Estimation and Predictive Momentum Management

### v1

@@ 4.4 Input Constraints
To ensure actuator feasibility and the satisfaction of hardware limits, constraints are imposed on both the control input magnitude and the rate of AMT motion. The actuator input vector $\mbf{u}_{j|t_{k}}$ is subject to the constraints $\mbf{u}_{\text{min}}\leq\mbf{u}_{j|t_{k}}\leq\mbf{u}_{\text{max}}$, where the bounds $\mbf{u}_{\text{min}}=\begin{bmatrix}\mbf{u}^{\text{AMT}^{\mathsf{T}}}_{\text{min}}&{u}^{\text{RCD}}_{\text{min}}\end{bmatrix}^{\mathsf{T}}$ and $\mbf{u}_{\text{max}}=\begin{bmatrix}\mbf{u}^{\text{AMT}^{\mathsf{T}}}_{\text{max}}&{u}^{\text{RCD}}_{\text{max}}\end{bmatrix}^{\mathsf{T}}$ reflect the physical actuation limits of the AMT and RCDs.
@@ 4.5 Objective Function
$\sum_{j=0}^{N-1}\Bigl(\mbf{x}_{j|t_{k}}^{\text{MPC}^{\mathsf{T}}}\mbf{Q}\mbf{x}^{\text{MPC}}_{j|t_{k}}+\mbf{u}_{j|t_{k}}^{\mathsf{T}}\mbf{R}\mbf{u}_{j|t_{k}}+\tilde{\mbf{u}}^{\text{AMT}^{\mathsf{T}}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}\Bigl)+\mbf{x}_{N|t_{k}}^{\text{MPC}^{\mathsf{T}}}\mbf{Q}_{N}\mbf{x}^{\text{MPC}}_{N|t_{k}}+{\boldsymbol{\alpha}}^{\mathsf{T}}\mbf{C}{\boldsymbol{\alpha}},$
where $\mbf{x}^{\text{MPC}}_{j|t_{k}}$ and $\mbf{u}_{j|t_{k}}=\begin{bmatrix}\mbf{u}^{\text{AMT}^{\mathsf{T}}}_{j|t_{k}}&{u}^{\text{RCD}}_{j|t_{k}}\end{bmatrix}^{\mathsf{T}}$ denote the predicted state and control input at stage $j$ over the prediction horizon of length $N$; $\mbf{Q}=\mbf{Q}^{\mathsf{T}}$ and $\mbf{R}=\mbf{R}^{\mathsf{T}}$ are positive semi-definite and positive definite weighting matrices, respectively, penalizing the state and control input; $\mbf{Q}_{N}$ is the terminal weighting matrix for the final predicted state at stage $N$; $\mbf{C}=\mbf{C}^{\mathsf{T}}$ is a positive semi-definite matrix that penalizes violation of the soft constraint via the slack variable ${\boldsymbol{\alpha}}\geq\mbf{0}$; ${\tilde{\mbf{R}}}={\tilde{\mbf{R}}}^{\mathsf{T}}$ is a positive semi-definite matrix penalizing the rate of AMT translation; and $\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}=\mbf{u}^{\text{AMT}}_{j|t_{k}}-\mbf{u}^{\text{AMT}}_{j-1|t_{k}}$ is the difference between the $j$-th AMT input and the previous input.
@@ 4.6 Summary of MPC Policy and Implementation Details
$\operatorname*{minimize}_{{\boldsymbol{\mathcal{X}}},\hskip 2.0pt{\boldsymbol{\mathcal{U}}},\hskip 2.0pt{\boldsymbol{\alpha}}}\sum_{j=0}^{N-1}\Bigl(\mbf{x}_{j|t_{k}}^{\text{MPC}^{\mathsf{T}}}\mbf{Q}\mbf{x}^{\text{MPC}}_{j|t_{k}}+\mbf{u}_{j|t_{k}}^{\mathsf{T}}\mbf{R}\mbf{u}_{j|t_{k}}+\tilde{\mbf{u}}^{\text{AMT}^{\mathsf{T}}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}\Bigl)+\mbf{x}_{N|t_{k}}^{\text{MPC}^{\mathsf{T}}}\mbf{Q}_{N}\mbf{x}_{N|t_{k}}^{\text{MPC}}+{\boldsymbol{\alpha}}^{\mathsf{T}}\mbf{C}{\boldsymbol{\alpha}}$ (13)
@@
$\hskip 10.00002pt\mbf{x}^{\text{MPC}}_{j+1|t_{k}}=\mbf{A}^{\text{MPC}}_{k}\mbf{x}^{\text{MPC}}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{w,k}\mbf{w}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{u1,k}\mbf{u}^{\text{AMT}}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{u2,k}{u}^{\text{RCD}}_{j|t_{k}},\qquad j=0,1,\ldots,N-1,$
@@
where ${\boldsymbol{\alpha}}\in\mathbb{R}^{4}$, ${\boldsymbol{\mathcal{X}}}=\{\mbf{x}^{\text{MPC}}_{0|t_{k}},\mbf{x}^{\text{MPC}}_{1|t_{k}},\ldots,\mbf{x}^{\text{MPC}}_{N|t_{k}}\}$, ${\boldsymbol{\mathcal{U}}}=\{\mbf{u}_{-1|t_{k}},\mbf{u}_{0|t_{k}},\mbf{u}_{1|t_{k}},\ldots,\mbf{u}_{N-1|t_{k}}\}$ are the design variables, $N$ is the number of timesteps in the prediction horizon, $\mbf{x}^{\text{MPC}}(t_{k})$ is the known system state at time $t_{k}$, and $\mbf{u}^{\text{AMT}}(t_{k})$ is the AMT position at time $t_{k}$.
Due to the use of a quadratic objective function, affine equality constraints, and affine inequality constraints, this MPC policy can be solved as a QP at each time step. Solving this problem yields a sequence of optimal control inputs over the prediction horizon, i.e., ${\boldsymbol{\mathcal{U}}}^{*}=\{\mbf{u}_{-1|t_{k}}^{*},\mbf{u}_{0|t_{k}}^{*},\mbf{u}_{1|t_{k}}^{*},\ldots,\mbf{u}_{N-1|t_{k}}^{*}\}$. Only the first input ($\mbf{u}_{0|t_{k}}^{*}$) is applied to the system before proceeding to the next time step and again solving for the optimal sequence of control inputs.
@@ 5.1 Simulation Setup
$\mbf{f}^{\text{SRP}}_{b}$ $\big[0.0003\,\,\,0\,\,\,0.013\big]^{\mathsf{T}}$ N
@@
[…]
Table 2: Kalman filter estimation tuning parameters used in the numerical simulations.
Parameter Value Unit
${\boldsymbol{\sigma}}_{\theta}$ $\text{diag}(1.86,1.86,12.5)\cdot 10^{-3}$ deg
${\boldsymbol{\sigma}}_{\omega}$ $\text{diag}(3.06,3.06,25)\cdot 10^{-4}$ deg/s
[…]
$\mbf{q}_{\theta}$ $0.01^{2}\cdot\mbf{1}_{3\times 3}$ deg2
$\mbf{q}_{\omega}$ $0.0001^{2}\cdot\mbf{1}_{3\times 3}$ (deg/s)2
[…]
$\mbf{R}^{\text{KF}}$ $\text{diag}({\boldsymbol{\sigma}}_{\theta}^{2},{\boldsymbol{\sigma}}_{\omega}^{2},{\boldsymbol{\sigma}}_{h}^{2},{\boldsymbol{\sigma}}_{e}^{2},{\boldsymbol{\sigma}}_{\dot{h}}^{2})$
$\mbf{Q}^{\text{KF}}_{\text{model}}$ $\text{diag}(\mbf{q}_{\theta},\mbf{q}_{\omega},\mbf{q}_{h},\mbf{q}_{e})$
$\mbf{Q}^{\text{KF}}_{\text{dist}}$ $\text{diag}({q}_{\tau 1},{q}_{\tau 2},{q}_{\tau 3})$
In this work, the Kalman filter operates at the same frequency as the momentum management system, which has a time step of $100$ seconds. The system undergoes a initial slew of attitude tracking, and the Kalman filter acquires its first measurement update at the first momentum management timestep, i.e., $t_{k}=100$ sec. After the measurement update, the momentum management policy determines the associated AMT and RCD inputs, which are then used in the time update using the Kalman filter process model. The momentum management inputs are applied to the nonlinear dynamics as in Eq. (1) until the next momentum management timestep, the process of a measurement update, momentum management input determination, time update, and application of the input to the nonlinear system is repeated.
@@ 5.2 NASA’s State-of-the-art Method
For practicality and for a fair comparison to the proposed method, the threshold-based momentum management uses state estimates from the Kalman filter to determine AMT and RCD inputs. Specifically, the angular momentum estimate $\hat{\mbf{h}}_{b,k}^{\text{RW}^{+}}$ is used to assess the activation/deactivation threshold and AMT proportional control, and $\hat{\dot{\mbf{h}}}_{b,k}^{\text{RW}^{+}}=\mbf{K}_{p}(\hat{{\boldsymbol{\theta}}}^{+}_{k}-{\boldsymbol{\theta}}_{d})+\mbf{K}_{d}(\hat{{\boldsymbol{\omega}}}^{{ba}^{+}}_{b,k}-\dot{{\boldsymbol{\theta}}}_{d})+\mbf{K}_{i}\hat{\mbf{e}}^{\text{int}^{+}}$ is used for the AMT derivative control. For the AMT integral control, it is assumed that a perfect measurement of $\int^{t}_{t_{0}}\mbf{h}^{\text{RW}}_{b}(\tau)\textrm{d}\tau$ is accessible in the ADCS.
@@
Figure 3: Simulation results using NASA’s Solar Cruiser momentum management strategy from Inness et al. (2023); Tyler et al. (2023), featuring RW saturation under a $4^{\circ}$ slew compared to a $3^{\circ}$ slew. The black dashed line in (c) denotes the activation threshold ($50\%$ of the soft constraint in MPC) , and the green dashed line denotes the deactivation threshold ($50\%$ of the activation threshold).
Using NASA’s state-of-the-art thresholding momentum management policy, a slew maneuver tracking initial attitude of ${\boldsymbol{\theta}}_{0}=\Big[-3^{\circ}\,\,\,0\,\,\,0\Big]^{\mathsf{T}}$ to ${{\boldsymbol{\theta}}}_{d}=\mbf{0}$ is performed. As shown in Fig. 3 with the label “NASA-3,” the $3^{\circ}$ slew demonstrates effective momentum management, where the performance is similar to that shown by Inness et al. (2023); Tyler et al. (2023), although this is difficult to compare quantitatively due to redacted plot axes. The momentum management method developed by Inness et al. (2023); Tyler et al. (2023) is effective at keeping the angular momentum of the RWs within reasonable bounds with realistic actuation inputs. However, with a slightly larger slew maneuver starting from ${\boldsymbol{\theta}}_{0}=\Big[-4^{\circ}\,\,\,0\,\,\,0\Big]^{\mathsf{T}}$, the system suffers from RW saturation, and the solar sail loses attitude control authority, which is shown in the result of Fig. 3 with the label “NASA-4.” For reference, the black dashed lines in Fig. 3(b) indicate $25\%$ of the angular momentum capacity of each RW, which is also the soft constraint value chosen for the proposed MPC-based approach in the following sections. The black dashed lines in Fig. 3(c) indicate the activation thresholds, while the green dashed lines represent the deactivation thresholds.

### v2

@@ 4.5 Input Constraints
To ensure actuator feasibility and the satisfaction of hardware limits, constraints are imposed on both the control input magnitude and the rate of AMT motion. The actuator input $\mbf{u}_{j|t_{k}}=\begin{bmatrix}\mbf{u}^{\text{AMT}^{\mathsf{T}}}_{j|t_{k}}&{u}^{\text{RCD}}_{j|t_{k}}\end{bmatrix}^{\mathsf{T}}$ is subject to the constraints $\mbf{u}_{\text{min}}\leq\mbf{u}_{j|t_{k}}\leq\mbf{u}_{\text{max}}$, where the bounds $\mbf{u}_{\text{min}}=\begin{bmatrix}\mbf{u}^{\text{AMT}^{\mathsf{T}}}_{\text{min}}&-\tau^{\text{RCD}}_{b3,\text{on}}\end{bmatrix}^{\mathsf{T}}$ and $\mbf{u}_{\text{max}}=\begin{bmatrix}\mbf{u}^{\text{AMT}^{\mathsf{T}}}_{\text{max}}&\tau^{\text{RCD}}_{b3,\text{on}}\end{bmatrix}^{\mathsf{T}}$ reflect the physical actuation limits of the AMT position and RCD torque. It is assumed that the RCD torque magnitude is accessible from the ADCS, which can be computed based on the SIA.
@@ 4.6 Objective Function
$\sum_{j=0}^{N-1}\Bigl((\mbf{x}_{j|t_{k}}^{\text{MPC}}-{\bar{\mbf{x}}}_{j|t_{k}}^{\text{MPC}})^{\mathsf{T}}\mbf{Q}(\mbf{x}^{\text{MPC}}_{j|t_{k}}-{\bar{\mbf{x}}}_{j|t_{k}}^{\text{MPC}})+\mbf{u}_{j|t_{k}}^{\mathsf{T}}\mbf{R}\mbf{u}_{j|t_{k}}+\tilde{\mbf{u}}^{\text{AMT}^{\mathsf{T}}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}\Bigl)+(\mbf{x}_{N|t_{k}}^{\text{MPC}}-{\bar{\mbf{x}}}_{N|t_{k}}^{\text{MPC}})^{\mathsf{T}}\mbf{Q}_{N}(\mbf{x}^{\text{MPC}}_{N|t_{k}}-{\bar{\mbf{x}}}_{N|t_{k}}^{\text{MPC}})+{\boldsymbol{\alpha}}^{\mathsf{T}}\mbf{C}{\boldsymbol{\alpha}},$
where $\mbf{x}^{\text{MPC}}_{j|t_{k}}$ and $\mbf{u}_{j|t_{k}}$ denote the predicted state and control input at stage $j$ over the prediction horizon of length $N$; $\bar{\mbf{x}}^{\text{MPC}}_{j|t_{k}}=\Big[{\boldsymbol{\theta}}_{d,j|t_{k}}^{\mathsf{T}}\,\,\,{\boldsymbol{\omega}}_{d,j|t_{k}}^{{\mathsf{T}}}\,\,\,\mathbf{M}_{34}^{\dagger}\hat{\mbf{h}}_{b,k}^{\text{RW}^{\mathsf{T}}}\,\,\,\hat{\mbf{e}}^{\text{int}^{\mathsf{T}}}_{k}\Big]^{\mathsf{T}}$ is the operating point at stage $j$ over the prediction horizon of length $N$; $\mbf{Q}=\mbf{Q}^{\mathsf{T}}$ and $\mbf{R}=\mbf{R}^{\mathsf{T}}$ are positive semi-definite and positive definite weighting matrices, respectively, penalizing the state and control input; $\mbf{Q}_{N}$ is the terminal weighting matrix for the final predicted state at stage $N$; $\mbf{C}=\mbf{C}^{\mathsf{T}}$ is a positive semi-definite matrix that penalizes violation of the soft constraint via the slack variable ${\boldsymbol{\alpha}}\geq\mbf{0}$; ${\tilde{\mbf{R}}}={\tilde{\mbf{R}}}^{\mathsf{T}}$ is a positive semi-definite matrix penalizing the rate of AMT translation; and $\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}=\mbf{u}^{\text{AMT}}_{j|t_{k}}-\mbf{u}^{\text{AMT}}_{j-1|t_{k}}$ is the difference between the $j$-th AMT input and the previous input.
@@ 4.7 Summary of MPC Policy and Implementation Details
$\operatorname*{minimize}_{{\boldsymbol{\mathcal{X}}},\hskip 2.0pt{\boldsymbol{\mathcal{U}}},\hskip 2.0pt{\boldsymbol{\alpha}}}\sum_{j=0}^{N-1}\Bigl((\mbf{x}_{j|t_{k}}^{\text{MPC}}-{\bar{\mbf{x}}}_{j|t_{k}}^{\text{MPC}})^{\mathsf{T}}\mbf{Q}(\mbf{x}^{\text{MPC}}_{j|t_{k}}-{\bar{\mbf{x}}}_{j|t_{k}}^{\text{MPC}})+\mbf{u}_{j|t_{k}}^{\mathsf{T}}\mbf{R}\mbf{u}_{j|t_{k}}+\tilde{\mbf{u}}^{\text{AMT}^{\mathsf{T}}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}\Bigl)+(\mbf{x}_{N|t_{k}}^{\text{MPC}}-{\bar{\mbf{x}}}_{N|t_{k}}^{\text{MPC}})^{\mathsf{T}}\mbf{Q}_{N}(\mbf{x}_{N|t_{k}}^{\text{MPC}}-{\bar{\mbf{x}}}_{N|t_{k}}^{\text{MPC}})+{\boldsymbol{\alpha}}^{\mathsf{T}}\mbf{C}{\boldsymbol{\alpha}}$ (14)
@@
$\hskip 10.00002pt\mbf{x}^{\text{MPC}}_{j+1|t_{k}}=\mbf{A}^{\text{MPC}}_{j|t_{k}}\mbf{x}^{\text{MPC}}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{w,j|t_{k}}\mbf{w}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{u1,j|t_{k}}\mbf{u}^{\text{AMT}}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{u2,j|t_{k}}{u}^{\text{RCD}}_{j|t_{k}}+\mbf{z}^{\text{MPC}}_{j|t_{k}},\qquad j=0,1,\ldots,N-1,$
@@
where ${\boldsymbol{\alpha}}\in\mathbb{R}^{4}$, ${\boldsymbol{\mathcal{X}}}=\{\mbf{x}^{\text{MPC}}_{0|t_{k}},\mbf{x}^{\text{MPC}}_{1|t_{k}},\ldots,\mbf{x}^{\text{MPC}}_{N|t_{k}}\}$, ${\boldsymbol{\mathcal{U}}}=\{\mbf{u}_{-1|t_{k}},\mbf{u}_{0|t_{k}},\mbf{u}_{1|t_{k}},\ldots,\mbf{u}_{N-1|t_{k}}\}$ are the design variables, $N$ is the number of timesteps in the prediction horizon, $\mbf{x}^{\text{MPC}}(t_{k})=\begin{bmatrix}\hat{{\boldsymbol{\theta}}}_{k}^{\mathsf{T}}&\hat{{\boldsymbol{\omega}}}^{ba^{\mathsf{T}}}_{b,k}&\mathbf{M}_{34}^{\dagger}\hat{\mbf{h}}_{b,k}^{\text{RW}^{\mathsf{T}}}&\hat{\mbf{e}}_{k}^{\text{int}^{\mathsf{T}}}\end{bmatrix}^{\mathsf{T}}$ is the known system state at time $t_{k}$, $\mbf{w}_{j|t_{k}}=\hat{\mbf{w}}^{+}_{k}$ is the Kalman filter disturbance estimate, and $\mbf{u}^{\text{AMT}}(t_{k})$ is the AMT position at time $t_{k}$.
Due to the use of a quadratic objective function, affine equality constraints, and affine inequality constraints, this MPC policy can be solved as a QP at each time step. The MATLAB function quadprog with its default settings is used to solve the QP optimization problem in the simulation results of this work. Solving this problem yields a sequence of optimal control inputs over the prediction horizon, i.e., ${\boldsymbol{\mathcal{U}}}^{*}=\{\mbf{u}_{-1|t_{k}}^{*},\mbf{u}_{0|t_{k}}^{*},\mbf{u}_{1|t_{k}}^{*},\ldots,\mbf{u}_{N-1|t_{k}}^{*}\}$. Only the first input ($\mbf{u}_{0|t_{k}}^{*}$) is applied to the system before proceeding to the next time step and again solving for the optimal sequence of control inputs.
@@ 5.1 Simulation Setup
The simulation parameters are chosen to reflect NASA’s Solar Cruiser (Johnson et al., 2019; Johnson and Curran, 2020; Tyler et al., 2023; Inness et al., 2023). The total mass of the sailcraft is $111$ kg, where the bus and the sail each contribute half of the total mass. The out-of-plane offset between the CM and the sail surface (also the CP) is captured by ${r}^{ps}_{b3}$, the third component of $\mbf{r}^{ps}_{b}(t)$. Unlike the simulations performed by Shen and Caverly (2026) that assumed this offset to be zero, this distance is chosen as ${r}^{ps}_{b3}=0.47$ m in this paper to more accurately simulate Solar Cruiser’s geometry. The rest of Solar Cruiser’s physical parameters are included in Table 1.
[…]
A static membrane shape model developed in Bunker and Caverly (2026) and motivated by the Solar Cruiser shape analysis performed by Gauvain and Tyler (2023) is employed to compute $\mbf{f}_{b}^{\text{SRP}}$ and ${\boldsymbol{\tau}}_{b}^{\text{dist}}$ based on the solar sail’s attitude $\boldsymbol{\theta}$}. This sail shape model implements Solar Cruiser’s membrane optical properties (Heaton and Artusio-Glimpse, 2015; Heaton et al., 2017), as well as its expected structural and membrane deformation (Gauvain and Tyler, 2023). A non-flat sail membrane with non-ideal reflectivity properties is used in the numerical simulations, where the billowing at the centroid of each quadrant is $15$ cm, and a $\pm 50$ cm alternating boom tip deflections to match the analysis of Bunker and Caverly (2026). The relationship between sailcraft attitude, SIA and clock angle, and the corresponding SRP force and disturbance torque at $1$ au is included in Table 2. Given the slow varying nature and small magnitude of SRP force and torque, $\mbf{f}^{\text{SRP}}_{b}$ and ${\boldsymbol{\tau}}^{\text{dist}}_{b}$ are updated at every $20$ seconds in the simulation. The simulation and attitude control timestep is $\textrm{d}t=1$ second, and the momentum management time step is $\Delta t=100$ seconds.
Table 2: Attitude correspondence to SRP force and torque.
Attitude $\boldsymbol{\theta}$} [SIA,clock angle] $\mbf{f}^{\text{SRP}}_{b}$ ${\boldsymbol{\tau}}^{\text{dist}}_{b}$
@@
The AMT and RCD actuation is as discussed in Section 2.4. The AMT has a translation limit of $\mbf{u}^{\text{AMT}}_{\text{max}}=-\mbf{u}^{\text{AMT}}_{\text{min}}=\Big[0.29\,\,\,0.29\Big]^{\mathsf{T}}$ m and a rate limit of $\dot{\mbf{u}}^{\text{AMT}}_{\text{max}}=-\dot{\mbf{u}}^{\text{AMT}}_{\text{min}}=\Big[0.5\,\,\,0.5\Big]^{\mathsf{T}}$ mm/s in the $\underrightarrow{b}^{1}$ and $\underrightarrow{b}^{2}$ axes (Johnson and Curran, 2020). The discrete-time rate constraint is defined as $\dot{\mbf{u}}^{\text{AMT}}_{\text{max}}=-\dot{\mbf{u}}^{\text{AMT}}_{\text{min}}=(\mbf{u}^{\text{AMT}}_{j+1|t_{k}}-\mbf{u}^{\text{AMT}}_{j|t_{k}})/\Delta t$, which limits the maximum AMT position change to be $0.05$ m in each axis at every momentum management time step. In the simulation ($\textrm{d}t=1$ second), the AMT is actuated to move at its maximum rate towards the momentum management commanded position (updated every $\Delta t=100$ seconds) in each of its axis until reaching the commanded position. The roll torque generated when the RCDs are turned on is set to meet the Solar Cruiser’s roll torque requirement as defined in Eq. (8).
@@
In this work, the Kalman filter operates at the same frequency as the momentum management system, which has a time step of $100$ seconds. The system undergoes an initial slew of attitude tracking, and the Kalman filter acquires its first measurement update at the first momentum management timestep, i.e., $t_{k}=100$ sec. After the measurement update, the momentum management policy determines the associated AMT and RCD inputs, which are then used in the time update using the Kalman filter process model. The momentum management input commands are passed through the AMT and RCD actuation dynamics as discussed in Section 2.4, and then applied to the nonlinear dynamics as in Eq. (1) until the next momentum management timestep. The process of a measurement update, momentum management input determination, time update, and application of the input to the nonlinear system is repeated.
@@ 5.2 NASA’s State-of-the-art Method
For practicality and for a fair comparison to the proposed method, the threshold-based momentum management uses state estimates from the Kalman filter to determine AMT and RCD inputs. Specifically, the angular momentum estimate $\hat{\mbf{h}}_{b,k}^{\text{RW}^{+}}$ is used to assess the activation/deactivation threshold and AMT proportional control, and $\hat{\dot{\mbf{h}}}_{b,k}^{\text{RW}^{+}}=\mbf{K}_{p}(\hat{{\boldsymbol{\theta}}}^{+}_{k}-{\boldsymbol{\theta}}_{d})+\mbf{K}_{d}(\hat{{\boldsymbol{\omega}}}^{{ba}^{+}}_{b,k}-{\boldsymbol{\omega}}_{d})+\mbf{K}_{i}\hat{\mbf{e}}^{\text{int}^{+}}_{k}$ is used for the AMT derivative control. For the AMT integral control, it is assumed that a perfect measurement of $\int^{t}_{t_{0}}\mbf{h}^{\text{RW}}_{b}(\tau)\textrm{d}\tau$ is accessible in the ADCS.
@@
Figure 3: Simulation results using NASA’s Solar Cruiser momentum management strategy from Inness et al. (2023); Tyler et al. (2023), featuring RW saturation under a slew of ${{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,10^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$ trajectory compared to a ${{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,10.5^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$ trajectory. The black dashed line in (c) denotes the activation threshold ($50\%$ of the soft constraint in MPC) , and the green dashed line denotes the deactivation threshold ($50\%$ of the activation threshold).
Using NASA’s state-of-the-art thresholding momentum management policy, a maneuver sequence tracking initial attitude of ${\boldsymbol{\theta}}_{0}=\mbf{0}$ to ${{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,10^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$ is performed, following the desired hold-slew-hold-slew-hold trajectory. As shown in Fig. 3 with the label “NASA,” the maneuver demonstrates effective momentum management, where the steady-state performance is similar to that shown by Inness et al. (2023); Tyler et al. (2023), although this is difficult to compare quantitatively due to redacted plot axes. The momentum management method developed by Inness et al. (2023); Tyler et al. (2023) is effective at keeping the angular momentum of the RWs within reasonable bounds with realistic actuation inputs. However, with a slightly larger slew maneuver with ${{\boldsymbol{\theta}}}_{\text{goal}}=\Big[0^{\circ}\,\,\,10.5^{\circ}\,\,\,1^{\circ}\Big]^{\mathsf{T}}$, the system suffers from RW saturation, and the solar sail loses attitude control authority, which is shown in the result of Fig. 3 with the label “NASA-sat.” For reference, the black dashed lines in Fig. 3(b) indicate $25\%$ of the angular momentum capacity of each RW, which is also the soft constraint value chosen for the proposed MPC-based approach in the following sections. The black dashed lines in Fig. 3(c) indicate the activation thresholds, while the green dashed lines represent the deactivation thresholds.